the normal to a parabola y^2=4ax at the point t,where t is not 0 meets the curve again at the point t'.find t' in terms of t.determine a point on the x-axis where the tangent at t' meets the x-axis.
1
Expert's answer
2021-07-30T07:16:23-0400

Parabola : "y^2 = 4ax"

Any point P on the parabola in parametric form is given as : "(at^2, 2at)"

The slope of tangent at any point is "\dfrac{1}{t}"


Hence, the slope of normal at point "(at^2, 2at)" is = "-t"


The equation of the normal becomes


"y - 2at = -t \ (x - at^2)"


"y = -t x + at^3 + 2at"


If the above equation intersects the curve again at point at ("at'^2, 2at'" )


Then it would satisfy the equation of normal. On substituting the value we get


"2at' = -att'^2 + at^3 + 2at"



"2a(t' - t) = at(t-t')(t+t')"


"2a = -at(t+t')"


"\dfrac{-2}{t} = t + t'"


"t' = \dfrac{-2}{t} - t"


Hence, the normal to the parabola meets the curve again at point "t' = \dfrac{-2}{t} - t"






Now, let us find the tangent to the curve at the point (at'2 , 2at' )


Slope of tangent at this point is "\dfrac{1}{t'}"


Equation of tangent is given by


"y - 2at' = \dfrac{1}{t'}(x - at'^2)"


To find the point where the above tangent meets the x-axis, we put y = 0.

Hence on substituting y = 0


"- 2at' = \dfrac{1}{t'}(x - at'^2)"


"x= -at'^2"


Hence, the tangent intersects the curve at ( "-at'^2,0" ).